OpenStudy (anonymous):

clearly a celebrity BWAHAHAHA;)now perform the indicated operation g(a)=3a-4 h(a)=3a squared-a.i know that to find (g-h)(a) you plug something in on the calculater RIGHT?

OpenStudy (anonymous):

No, you subtract the two equations.

OpenStudy (anonymous):

if \(g = 3a-4\) and \(h = 3a^2 -a\), what is g - h?

OpenStudy (anonymous):

(g-h)(a)=g(a)-h(a)

OpenStudy (anonymous):

ok i get that u subtract but what do u do with the (a)?

OpenStudy (anonymous):

The (a) is just notation that tells you that g and h depend on a.